A radio-controlled model airplane has a momentum given by [(βˆ’0.75π‘˜π‘” βˆ™ π‘š 𝑠 3 ⁄ )𝑑 2 + (3.0π‘˜π‘” βˆ™ π‘šβ„π‘ )]𝑖̂+ (0.25π‘˜π‘” βˆ™ π‘š 𝑠 2 ⁄ )𝑑𝑗̂. What are the π‘₯-, 𝑦-, and 𝑧-components of the net force on the airplane?


1
Expert's answer
2021-12-14T15:09:17-0500

"\vec p=(-0.75t^2+3)\cdot\vec i+0.25t\cdot\vec j"


"\vec F=d\vec p/dt=-1.5t\cdot \vec i+0.25\cdot\vec j"


"F_x=-1.5t" , "F_y=0.25" , "F_z=0"